OpenStudy (anonymous):

dy/dx = 7/x how do you separate this????

OpenStudy (anonymous):

what do you mean by separate it?

OpenStudy (anonymous):

separable differential equations!!

OpenStudy (anonymous):

if \(\frac{dy}{dx}=\frac{7}{x}\) then \(y=7\ln(x)\)

OpenStudy (anonymous):

ohh ok thanks :)